OpenStudy (alphadxg):

Phenylketonuria (PKU) is an autosomal recessive human disorder. Individuals with PKU have a defective phenylalanine hydroxylase gene (called the “H” gene in this problem). Individuals with PKU are at risk for mental retardation unless phenylalanine is restricted from their diet. Given the importance of diet for individuals with this disorder, newborns in the U.S. are screened for PKU. The chart below shows the frequency of PKU genotypes in the U.S. Caucasian population.

OpenStudy (alphadxg):

Assuming that this population is at Hardy-Weinberg equilibrium, what is the expected frequency of individuals who are homozygous dominant in the next generation? GENOTYPE Number of individuals out of 100 with this genotype HH 90 Hh 8 hh 2 Answer a. 0.02 b. 0.18 c. 0.884 d. 0.9 Last question I need.

OpenStudy (alphadxg):

@aaronq Last one I need I want to make sure with you that I have the write answer, I think I know what it is.

OpenStudy (aaronq):

hmm i vaguely remember doing this type..so they're asking for HH, which is \(p^2\).. so \(f(H)=\dfrac{188}{200}=0.94=p\) so \(p^2=0.94^2=0.8836\)